### v2

# Forensics Adapter: Adapting CLIP for Generalizable Face Forgery Detection

Abstract
We describe the Forensics Adapter, an adapter network designed to transform CLIP into an effective and generalizable face forgery detector. Although CLIP is highly versatile, adapting it for face forgery detection is non-trivial as forgery-related knowledge is entangled with a wide range of unrelated knowledge. Existing methods treat CLIP merely as a feature extractor, lacking task-specific adaptation, which limits their effectiveness. To address this, we introduce an adapter to learn face forgery traces – the blending boundaries unique to forged faces, guided by task-specific objectives. Then we enhance the CLIP visual tokens with a dedicated interaction strategy that communicates knowledge across CLIP and the adapter. Since the adapter is alongside CLIP, its versatility is highly retained, naturally ensuring strong generalizability in face forgery detection. With only trainable parameters, our method achieves a significant performance boost, improving by approximately on average across five standard datasets. We believe the proposed method can serve as a baseline for future CLIP-based face forgery detection methods. The code is available at https://github.com/OUC-VAS/ForensicsAdapter.
1 Introduction
Face forgery techniques111Face forgery strictly refers to techniques that manipulate local facial content, such as altering lip movement, swapping central face, editing expression, etc. Whole facial image synthesis using GAN or Diffusion models are out of this scope. have seen remarkable progress in recent years, largely due to significant advancements in generative models [18, 28, 23, 55]. These techniques enable the manipulation of faces with high realism, causing serious concerns to social security, such as privacy violations [7, 34], economic fraud [5, 31], etc. These concerns have motivated a great need for exploration into face forgery detection.
Recent face forgery detection methods are typically developed on CNNs, benefiting from the learning capacity of these models. They leverage various forgery clues, including biological signals [10, 72, 52], blending artifacts [50, 38, 48], and frequency signals [53, 47, 43]. While these methods can achieve favorable and near-perfect performance on standard datasets, their ability to detect unseen forgeries drops significantly. This is because these methods tend to overfit the specific forgeries present in the training data, limiting their generalization across different distributions. With new forgery techniques constantly emerging in real-world scenarios, these detection methods remain highly constrained in practical use. To address this issue, several generalizable face forgery detection methods have been proposed, aiming to capture common forgery clues through approaches such as pseudo-fake face augmentations [33, 57], and disentanglement learning [69, 42]. Nevertheless, given the limited variety of forgery categories, these methods still struggle to acquire sufficient forgery priors.
CLIP [54], a recent large vision-language model, contains extensive knowledge priors, including that related to forgery, offering promising potential to enhance generalizable face forgery detection. Inspired by this, several methods [41, 26, 29] are proposed to utilize CLIP for face forgery detection. They commonly use the frozen CLIP image encoder as a feature extractor, with an optional adapter network appended after CLIP. To learn forgery-related knowledge, the adapter network is trained using general objectives (e.g., cross-entropy) on face forgery datasets (see Fig. 1 (a)).
However, this straightforward adaptation has two major limitations: (1) Task-agnostic adaptation. The adapter is designed generically, rather than tailored to capture unique task-specific clues in face forgery. (2) Lack of interaction between CLIP and the adapter. Since CLIP is not specifically designed for this task, its forgery-related knowledge is intertwined with a broad range of other knowledge. The adapter, therefore, should interact more directly with CLIP to guide it toward forgery-related knowledge, rather than only refining the final output of CLIP. Moreover, the versatility of CLIP can serve as a powerful source of instruction for training the adapter effectively.
In this paper, we propose Forensics Adapter, a devoted framework for adapting CLIP to generalizable face forgery detection (see Fig. 1 (b)). Unlike previous CLIP-based methods, our method is grounded in the core nature of face forgery, with a task-specific adaptation that focuses on key forgery-related clues. Our motivation stems from the observation that face forgeries often involve localized manipulations, where synthesized content is blended back into the original regions, creating inconsistencies between the manipulated and original regions. These inconsistencies form blending boundaries, which serve as crucial clues specific to face forgery. To capture these subtle traces, we propose a lightweight adapter network alongside CLIP and introduce task-specific objectives to instruct learning of the adapter. Then we describe the interaction strategy between the adapter and CLIP, assisting the learning of the adapter while enhancing the token effectiveness of CLIP. With this adapter, CLIP is transformed into an effective face forgery detector.
While the blended inconsistencies have been utilized in previous methods [49, 57], the solution of specific adaptation to CLIP is novel and unexplored, offering fresh insights for future research. Extensive experiments demonstrate that our method outperforms state-of-the-arts by a large margin ( in AUC) across five public datasets, with only trainable parameters. We believe our method can establish a strong baseline for CLIP-based face forgery detection methods.
Our contribution can summarized in threefold: (1) We bridge the gap between CLIP and face forgery task by introducing a Forensics Adapter that effectively adapts CLIP for generalizable face forgery detection. (2) We provide a task-specific adapter design, including both architecture and objectives, and outline an interaction strategy between CLIP and the adapter. (3) The experimental results are promising, demonstrating the potential of our method as a baseline for CLIP-based face forgery detection methods.
2 Related Works
Face Forgery. Face forgery involves manipulating the local content of the original faces using generative strategies or 3D-based strategies [56, 40, 16, 15, 12]. Its pipeline usually consists of two steps. The first step is to create fake content, e.g., tampered lip or head movement, forged facial identity, etc. The second step is to blend the fake content back to the corresponding position of the original faces while retaining high visual quality. Although the recent face forgery techniques have greatly evolved, this two-step pipeline always introduces inconsistencies between the forgery and original areas, resulting in a blending boundary.
Besides face forgery, whole facial image synthesis is another widely studied forensics topic. It involves synthesizing whole images from scratch using GAN [18] or Diffusion models [23, 55], based on random noises or certain conditioned inputs. It is important to clarify that this paper focuses on face forgery rather than whole facial image synthesis.
Face Forgery Detection. Detecting face forgeries is a significant topic in recent years [39, 53, 49, 11, 3, 62]. With the great advancement of deep learning, recent methods are typically designed on Deep Neural Networks (DNNs) to capture the forgery traces, utilizing features ranging from the biological signals (e.g., eye blinking [39], heartbeat rhythm [52], head pose [72], facial action units [1]), frequency signals (e.g., [47, 53, 43]), and auto-learned clues with dedicated training schemes (e.g., [58, 6, 11]). However, these methods usually rely on specific training datasets, which can lead to overfitting on known data distributions and hinder their ability to generalize across different datasets. To enhance detection generalization, many methods have been proposed to create pseudo-fake faces by simulating blending boundaries in real images [37, 57, 33, 49, 3]. While these methods improve the generalization to a certain extent, limited data diversity still restricts detectors from learning generic features.
More recently, several efforts have started adapting CLIP [54] for face forgery detection [41, 26, 32]. For instance, FFAA [26] creates a dataset comprising real and forged faces with descriptions and forgery reasoning and implements the face forgery through a multi-answer decision strategy based on CLIP. GM-DF [32] attempts to mitigate discrepancies across multiple combined training datasets by leveraging CLIP. FCG [21] leverages CLIP to guide the detection of key facial components over temporal sequences. VLFFD [59] generates mixed forgery images with fine-grained prompts and uses them with original data for joint training. RepDFD [41] reprograms CLIP by merging universal perturbations to the visual input, without adjusting the inner parameters of the VLM model. CLIPping [29] explores the effectiveness of CLIP in combination with recent adaptation methods for universal deepfake detection.
Regardless of how advanced these methods are, they share a common limitation: they solely use CLIP as a feature extractor, overlooking the unique characteristics of face forgery detection. This results in limited task-specific adaptation and incomplete utilization of CLIP’s versatility. Therefore, we describe the Forensics Adapter, designed to capture specific face forgery traces and effectively adapt CLIP into a generalizable face forgery detector.
3 Forensics Adapter
To ensure practical usability, the adapter should be lightweight. Thus, we employ the tiny vision transformer [65] as the foundation of the adapter, incorporating only the first eight layers and positioning it parallel to CLIP to unleash its full capacity, inspired by the methods in [44, 66]. Unlike general computer vision tasks, the adapter is instructed to detect the blending boundaries in forged faces, with dedicated objectives (Sec. 3.1). Meanwhile, we describe an interaction strategy between the adapter and CLIP. On one hand, absorbing CLIP visual tokens can support the learning process of the adapter, while on the other, the task-specific knowledge gained by the adapter can be transferred back into CLIP (Sec. 3.2). With these task-specific objectives and interaction strategies, this adapter can effectively instruct CLIP to focus more on forgery traces while preserving its generalizability. The pipeline of our method is shown in Fig. 2.
3.1 Learning Task-specific Traces
Input Tokens. Given an input face image, we first divide it into patches and then convert them as visual tokens. These tokens are concatenated with learnable query tokens as the input for the adapter.
Prediction Head. To instruct the adapter in learning blending boundaries, we design a prediction head in the adapter to predict the blending boundary map. The input of this head is query tokens and visual tokens from the adapter respectively. For query tokens, we project it into -dimension using MLPs, which is denoted as . Similarly, we project visual tokens into -dimension using another MLPs, which is denoted as , where is of input image size. The inner product of and can generate the blending boundary map as
where denotes a couple of convolution operations that gradually transform the dimension of the output as . Each element in represents the probability of being on a blending boundary. To supervise the prediction and compel CLIP with the adapter to more accurately capture the subtle distinctions between real and fake regions, we generate ground truth blending boundaries for each input and propose three objectives regarding masked blending boundary learning, inter-face contrastive learning, and intra-face patch-wise contrastive learning. The details are elaborated below.
Masked Blending Boundary Learning. Following previous works [37], we first perform Gaussian blurring on ground truth manipulation mask and then generate ground truth blending boundary with . A straightforward way would be to directly calculate the Mean Square Error (MSE) between and . However, since the boundary occupies only a small portion of the map, using standard MSE can easily be influenced by the non-boundary areas, leading to suboptimal results. Therefore, we introduce a masking strategy to highlight the effect of the boundary areas. Specifically, we define a binary mask and check all patches of . If no boundary exists in a patch, the corresponding elements in are set to . Otherwise, they are set to 1. This objective can be defined as
Patch-wise Contrastive Learning. Note that in a forged face, the representations of the forged region should differ from the authentic regions. Highlighting this difference can assist the model in capturing the task-specific traces. To achieve this, we describe a patch-wise contrastive learning loss performed on the adapter and CLIP.
For the adapter, we extract the intermediate features and perform contrastive learning on them. Each element corresponds to a divided patch, which is labeled as fake if the forgery region covers more than of the corresponding patch, and as real otherwise. Denote as a feature, where are two sets containing real and fake elements. The patch-wise contrastive learning can be formulated as
where are elements from same set, represents the opposite set of , i.e., if , and otherwise. is the temperature parameter and is the cosine similarity.
For CLIP, we apply the same formulation to refine its visual tokens. Note that the parameters of CLIP are frozen and cannot be directly updated. Therefore, we employ another trainable convolution on the visual tokens (e.g., ) and perform patch-wise contrastive learning on the newly obtained tokens (e.g., ). Then we add the enhanced tokens back to CLIP with a factor , that is .
Sample-wise Contrastive Learning. In addition to the patch-wise correlation, there are certain associations between real and fake samples. Since the fake samples in this context may belong to different categories (e.g., DF, NT, etc.), we do not consider the similarity among them. Instead, we only focus on pulling close real samples while increasing the distance between real and fake samples.
Since CLIP focuses more on the global view, we perform sample-wise contrastive learning on it. Denote a training batch as , where denote the sets with real and fake faces. This learning process can be written as
where are samples from same set, represents the opposite set of as in Eq. (3). Since CLIP is frozen, we use the same strategy in patch-wise contrastive learning to update CLIP. Similarly, we employ another trainable convolution on the visual tokens (e.g., ) and perform the learning on the newly obtained tokens. Then we add the enhanced tokens back to CLIP with a the same factor .
3.2 Interaction between Adapter and CLIP
Absorbing CLIP Visual Tokens. The features of CLIP are highly versatile and can greatly support the learning process of the adapter. Thus we fuse these features into the adapter. Since CLIP and the adapter have different goals, we limit the feature fusion to the shallow layers of the adapter, avoiding interference with the learning of task-specific features. For a standard 24-layer ViT-L/14 model, we fuse features from layers to the layers of the adapter. Since the feature dimensions of CLIP and the adapter are inconsistent, we employ a convolution to align them. We then add these two features as the fused ones.
Enhancing CLIP Visual Tokens. To make CLIP focus more on blending boundary traces, we convey the knowledge of the adapter to CLIP without modifying any parameters. We achieve this by utilizing an attention bias strategy similar to the one proposed in [66]. Note that CLIP tokens comprise visual tokens and token, denoted as . For each layer, we duplicate copies of token to form an independent token as . The entire tokens at a layer can be written as . Then we update using self-attention based on the output of the adapter and original visual tokens in CLIP, as
where represents the layer number, denotes the query of , denotes the key of visual tokens , and indicates the value of visual tokens , respectively. Notably, is the attention bias generated by the end of the adapter, following similar operations in generating the blending boundary map in Eq. (1). Specifically, we use another MLPs to project visual tokens into -dimension as and generate the attention bias using the inner product of and as
where denotes the same query tokens in the prediction head. This attention bias adjusts the self-attention calculation in CLIP, serving as the bridge to convey knowledge of the adapter. Finally, we use for identification of face authenticity.
3.3 Overall Objectives
In the training phase, the main objective is a Cross-entropy loss applied on in the last layer of CLIP, denoted as , to achieve forgery detection. Together with the aforementioned objectives in Sec. 3.1, the overall objective can be written as
where are weighting factors to balance each term.
| Method | Venue | CDF-v1 | CDF-v2 | DFDC | DFDCP | DFD | Avg. |
| Xception [56] | ICCV’19 | 0.779 | 0.737 | 0.708 | 0.737 | 0.816 | 0.755 |
| EfficientB4 [60] | ICML’19 | 0.791 | 0.749 | 0.696 | 0.728 | 0.815 | 0.756 |
| F3Net [53] | AAAI’20 | 0.777 | 0.735 | 0.702 | 0.735 | 0.798 | 0.749 |
| X-ray [37] | CVPR’20 | 0.709 | 0.679 | 0.633 | 0.694 | 0.766 | 0.696 |
| FFD [11] | CVPR’20 | 0.784 | 0.744 | 0.703 | 0.743 | 0.802 | 0.755 |
| SPSL [43] | CVPR’21 | 0.815 | 0.765 | 0.704 | 0.741 | 0.812 | 0.767 |
| SRM [47] | CVPR’21 | 0.793 | 0.755 | 0.700 | 0.741 | 0.812 | 0.760 |
| Recce [6] | CVPR’22 | 0.768 | 0.732 | 0.713 | 0.734 | 0.812 | 0.752 |
| SBI [57] | CVPR’22 | - | 0.813 | - | 0.799 | 0.774 | - |
| UCF [69] | ICCV’23 | 0.779 | 0.753 | 0.719 | 0.759 | 0.807 | 0.763 |
| ED [2] | AAAI’24 | 0.818 | 0.864 | 0.721 | 0.851 | - | - |
| LSDA [71] | CVPR’24 | 0.867 | 0.830 | 0.736 | 0.815 | 0.880 | 0.826 |
| CFM [46] | TIFS’24 | - | 0.828 | - | 0.758 | 0.915 | - |
| Ours | - | 0.914 | 0.900 | 0.843 | 0.890 | 0.933 | 0.896 |
4 Experiments
4.1 Experimental Settings
Datasets. Following previous methods [71, 70], we utilize the following six publicly available datasets: FaceForensics++ (FF++) [56], Deepfake Detection Challenge (DFDC) [16], preview version of DFDC (DFDCP) [15], two versions of CelebDF (CDF-v1, CDF-v2) [40], and DeepfakeDetection (DFD) [12], respectively. To validate the generalizability, our method is trained on c23 compression version of FF++ and tested on other datasets.
| Method | Venue | CDF-v2 | DFDC | DFDCP |
| RealForensics [20] | CVPR’22 | 0.869 | 0.759 | - |
| TALL [67] | ICCV’23 | 0.908 | 0.768 | - |
| AltFreezing [64] | CVPR’23 | 0.895 | - | - |
| SeeABLE [33] | ICCV’23 | 0.873 | 0.759 | 0.863 |
| IID [25] | CVPR’23 | 0.838 | - | 0.812 |
| TALL++ [68] | IJCV’24 | 0.920 | 0.785 | - |
| SAM [8] | CVPR’24 | 0.890 | - | - |
| SBI [57] | CVPR’22 | 0.932 | 0.724 | 0.862 |
| AUNet [3] | CVPR’23 | 0.928 | 0.738 | 0.862 |
| CADDM [17] | CVPR’23 | 0.939 | 0.739 | - |
| SFDG [63] | CVPR’23 | 0.758 | 0.736 | - |
| LAA-NET [49] | CVPR’24 | 0.954 | - | 0.869 |
| LSDA [71] | CVPR’24 | 0.911 | 0.770 | - |
| CFM [46] | TIFS’24 | 0.897 | - | 0.802 |
| Ours | - | 0.957 | 0.872 | 0.929 |
Implementation Details. Our method is implemented by Pytorch 2.3.0 [51] with one Nvidia GTX 3090 GPU. The adapter is configured with a patch size of 16, an image resolution of 256, and no pre-trained weights are loaded. CLIP architecture is ViT-L/14 with a patch size of 14 and an image resolution of 224, with pre-trained weights loaded. The dimension is set to 128. For the adapter, we use the features from intermediate layers 4, 5, and 6 for patch-wise contrastive learning. For CLIP, we add a trainable convolution to its 13th layer to support both sample-wise contrastive learning and patch-wise contrastive learning, with the factor set to 0.05. For overall objectives, is different on the adapter and CLIP: on the adapter and on CLIP. Other factors are set to respectively. The entire model was trained on the FF++ c23 training set, with a batch size of 16, using the Adam optimizer [30] with a learning rate of 0.0002 and weight decay of 0.0005. We follow the configuration of DeepfakeBench [70]: For each video, 32 frames were extracted for training or testing, and RetinaFace [13] was used for face extraction.
4.2 Results
Compared with State-of-the-arts. For a comprehensive comparison, we compare our method with other state-of-the-arts under both frame-level and video-level evaluations.
Table 1 shows the cross-dataset evaluation results in frame-level AUC. In this scenario, all methods are trained on the FF++ c23 set and evaluated across different datasets. Note that the results of CFM [46] and ED [2] are taken from their original papers. Since SBI [57] does not report its frame-level results, we use the reproduced results from CFM. For all the rest methods, we take the results from the second-best method LSDA [71]. It can be observed that our method outperforms the previous best method LSDA by an impressive improvement in average AUC. Notably, on the widely recognized challenging dataset DFDC, our method significantly surpasses LSDA by in AUC, demonstrating the generalizability of our method.
Table 2 shows the cross-dataset evaluation results in video-level AUC. Following previous works [33], the evaluation is performed on CDF-v2, DFDC, and DFDCP datasets. The top section presents video-based methods and the bottom section corresponds to frame-based methods222Since several frame-based methods only report either frame-level AUC or video-level AUC, the included methods in Table 1 and Table 2 (bottom section) are not fully consistent.. Note that RealForensics [20], SeeABLE [33], SBI [57], AUNet [3] and LAA-NET [49] are trained on the custom dataset, while others are trained on FF++ c23 set. We can observe that our method significantly outperforms others on all datasets, improving AUC by on CDF-v2, DFDC, and DFDCP, respectively. These results also corroborate the effectiveness of our method in generalizable face forgery detection.
Compared with CLIP-based Methods. To fully demonstrate the effectiveness of the adapter, we compare our method with CLIP-based methods. Table 3 shows the comparison results of our method and other CLIP-based methods. Note that VLFFD [59], CLIPping [29], and FFAA [26] are trained on the custom dataset, while the others are trained on FF++ c23 set. Vanilla CLIP refers to the frozen ViT-L/14 model with an additional FC head. We can see that our method notably outperforms others on all datasets, improving the performance by on CDF-v2, DFDC, and DFD respectively. It further demonstrates the superiority of the proposed adapter.
| Method | Venue | CDF-v2 | DFDC | DFD |
| Vanilla CLIP [54] | ICML’21 | 0.777 | 0.742 | 0.834 |
| VLFFD [59] | arXiv’23 | 0.848 | - | 0.948 |
| FFAA [26] | arXiv’24 | - | 0.740 | 0.920 |
| GM-DF [32] | arXiv’24 | 0.832 | 0.772 | - |
| FCG [21] | arXiv’24 | 0.923 | 0.812 | - |
| RepDFD [41] | arXiv’24 | 0.899 | 0.810 | - |
| CLIPping [29] | ICMR’24 | - | 0.719 | 0.866 |
| Ours | - | 0.957 | 0.872 | 0.972 |
4.3 Analysis
| CDF-v1 | CDF-v2 | DFDC | DFDCP | DFD | |||||||||||||
| AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | |||
| 0.887 | 0.925 | 19.3 | 0.860 | 0.926 | 22.3 | 0.825 | 0.857 | 25.5 | 0.848 | 0.921 | 23.5 | 0.918 | 0.990 | 15.1 | |||
| 0.818 | 0.869 | 26.1 | 0.822 | 0.885 | 25.8 | 0.790 | 0.805 | 27.8 | 0.826 | 0.896 | 23.9 | 0.902 | 0.987 | 16.9 | |||
| 0.904 | 0.934 | 18.1 | 0.899 | 0.941 | 18.8 | 0.840 | 0.871 | 24.5 | 0.893 | 0.943 | 19.0 | 0.920 | 0.990 | 15.2 | |||
| 0.914 | 0.940 | 15.8 | 0.900 | 0.945 | 18.2 | 0.843 | 0.873 | 23.9 | 0.890 | 0.942 | 19.7 | 0.933 | 0.992 | 13.4 | |||
| ADP | CLIP | CDF-v1 | CDF-v2 | DFDC | DFDCP | DFD | ||||||||||
| AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | ||
| 0.814 | 0.895 | 24.9 | 0.789 | 0.880 | 28.7 | 0.744 | 0.778 | 32.3 | 0.748 | 0.868 | 33.0 | 0.854 | 0.982 | 22.4 | ||
| 0.845 | 0.912 | 24.0 | 0.822 | 0.899 | 25.8 | 0.778 | 0.817 | 29.8 | 0.809 | 0.893 | 28.3 | 0.899 | 0.987 | 18.5 | ||
| 0.881 | 0.926 | 20.3 | 0.875 | 0.923 | 21.4 | 0.817 | 0.840 | 26.2 | 0.833 | 0.894 | 24.0 | 0.903 | 0.988 | 17.3 | ||
| 0.914 | 0.940 | 15.8 | 0.900 | 0.945 | 18.2 | 0.843 | 0.873 | 23.9 | 0.890 | 0.942 | 19.7 | 0.933 | 0.992 | 13.4 | ||
Effect of Losses. This part studies the effect of each loss (, , ). We employ Area Under Curve (AUC), Average Precision (AP), and Equal Error Rate (EER) for evaluation. The results are shown in Table 4. It can be seen that using all losses achieves the best performance. Notably, seemingly has the largest impact on the performance, which averagely improves in AUC and AP , respectively, while reducing the EER by an average of .
Where Using Contrastive Learning. Theoretically, patch-wise contrastive learning and sample-wise contrastive learning can be used on both CLIP and adapter. Thus, we study where the learning should added. The results, shown in Table 5, highlight the importance of the proposed task-specific learning strategies. By using the learning on the adapter (ADP) side, we can improve the performance by in AUC and AP while reducing EER by on average. By applying the learning solely on the CLIP side, we can achieve a performance increase by around in AUC and AP while lowering EER by on average. When adding the learning strategy to both CLIP and the adapter, our method achieves the best performance.
Using Various LLMs. This part investigates the effect of using various LLMs. Specifically, we validate three LLMs: BLIP ViT-L/14 [36], CLIP ViT-B/16 [54], and CLIP ViT-L/14 [54] (the one used in main experiment). The results are shown in Table 6. We observe that BLIP is not suitable for the task, which highly degrades the performance compared to CLIP variants. This is possible because BLIP is trained using only 129 million image-text pairs and designed for image-text retrieval and VQA tasks with a bootstrapping mechanism. In contrast, CLIP is trained on a much larger dataset containing 400 million samples, enabling superior zero-shot capabilities in the modality of images. Straightforwardly, using large ViT can obtain more performance gain than the base version, which aligns with our understanding.
| LLMs | CDF-v2 | DFDC | DFDCP |
| BLIP ViT-L/14 [36] | 0.607 | 0.539 | 0.548 |
| CLIP ViT-B/16 [54] | 0.837 | 0.775 | 0.799 |
| CLIP ViT-L/14 [54] | 0.900 | 0.843 | 0.890 |
Using Various Fine-Tuning Methods. We also study the effect of using various fine-tuning methods, including LoRA [24] and subset fine-tuning [22]. As shown in Table 7, without task-specific adaptation, other fine-tuning methods yield only moderate performance.
| Method | CDF-v2 | DFDCP |
| LoRA [24] | 0.797 | 0.800 |
| subset fine-tuning [22] | 0.770 | 0.807 |
| Ours | 0.900 | 0.890 |
Using Various Adapter Architectures. In this section, we explore the effect of different adapter architectures, including ViT-tiny (ours), ViT-small, ViT-base, and ViT-large. All adapters are trained using the same configuration as in our main experiment. The results are presented in Table 8. As shown, performance does not improve as the number of trainable parameters increases under the same training configuration. This is likely because, with limited training resources, the smaller ViT-tiny architecture can be fully trained to capture forgery traces, whereas the larger architectures do not show additional benefits.
| Adapter | CDF-v1 | CDF-v2 | DFDC | DFDCP | DFD | Avg. |
| Large | 0.900 | 0.891 | 0.832 | 0.840 | 0.905 | 0.874 |
| Base | 0.895 | 0.869 | 0.833 | 0.862 | 0.902 | 0.872 |
| Small | 0.897 | 0.886 | 0.825 | 0.861 | 0.906 | 0.875 |
| Tiny | 0.914 | 0.900 | 0.843 | 0.890 | 0.933 | 0.896 |
Exploration of the Text Modality. In addition to using the visual encoder of CLIP, we also explore the incorporation of text encoder. Following CoOP [74], we design learnable contexts within the prompts and combine the text tokens with visual tokens. Table 9 presents the results. Surprisingly, the performance after using text modality is notably lower than only using the visual modality. This may be because the forged faces are mainly reliant on visual representations, and the integration of text modality introduces irrelevant noise, which distracts attention from the correct forgery traces.
| Modality | CDF-v1 | CDF-v2 | DFDC | DFDCP | DFD | Avg. |
| T+V | 0.738 | 0.720 | 0.726 | 0.719 | 0.814 | 0.744 |
| V | 0.914 | 0.900 | 0.843 | 0.890 | 0.933 | 0.896 |
| Method | StarGAN [9] | CramerGAN [4] | MMDGAN [35] | ADM [14] | PNDM [45] | ||||||||||
| AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | AUC | AP | EER | |
| IID [25] | 0.699 | 0.697 | 34.7 | 0.639 | 0.617 | 40.3 | 0.565 | 0.552 | 45.3 | 0.609 | 0.617 | 42.0 | 0.279 | 0.391 | 63.3 |
| LSDA [71] | 0.772 | 0.796 | 29.0 | 0.675 | 0.646 | 35.8 | 0.621 | 0.592 | 40.3 | 0.577 | 0.549 | 42.8 | 0.572 | 0.546 | 40.6 |
| Ours | 0.954 | 0.950 | 10.9 | 0.936 | 0.945 | 13.7 | 0.896 | 0.911 | 18.6 | 0.806 | 0.793 | 26.3 | 0.638 | 0.624 | 39.3 |
Detecting Whole Face Image Synthesis. Although our method is restricted to detecting face forgeries, we explore its potential for detecting whole face images synthesized by GANs and Diffusion models. Specifically, we select three GANs, including StarGAN [9], CramerGAN [4], and MMDGAN [35], and two Diffusion models, including ADM [14] and PNDM [45]. Table 10 shows the performance of our method compared to IID and LSDA. The results indicate that, despite being designed specifically for face forgery tasks, our method exhibits satisfied performance, and still surpasses the counterparts by a large margin on both GANs and Diffusion models.
Robustness. This part studies the performance of our method confronting various perturbations. We employ the same perturbation configurations following [27], which includes five levels of six different types of perturbations: color saturation, color contrast, block wise, gaussian blur, jpeg compression, gaussian noise. We compare our method with LSDA, IID, and vanilla CLIP methods in video-level AUC. The results are shown in Fig. 3. It can be seen that our method is more robust than others. All methods are trained on the FF++ c23 dataset.
Complexity Analysis. Since we freeze the CLIP model and only train the lightweight adapter, we significantly reduce the training overhead to 5.7M parameters. Table 11 compares the parameter counts and video-level AUC performance of various methods, all trained on the FF++ c23 dataset. Notably, other methods require three to six times more trainable parameters than ours, yet our method achieves a substantial performance improvement, highlighting the effectiveness of our method.
| Method | # Params | CDF-v2 | DFDC |
| LipForensics [19] | 36.0M | 0.824 | 0.735 |
| FTCN [73] | 26.6M | 0.869 | 0.740 |
| DCL [58] | 19.35M | 0.823 | 0.767 |
| CFM [46] | 25.37M | 0.897 | 0.802 |
| Ours | 5.7M | 0.957 | 0.872 |
T-SNE Visualization. To demonstrate our adapter greatly improves the effectiveness of CLIP, we visualize the output of CLIP before and after using the adapter on the CDF-v2 dataset. Specifically, we randomly select 500 real and 500 fake faces from the testing set and employ the T-SNE algorithm [61] for visualization. The results, shown in Fig. 4, reveal that the features of real and fake faces are highly separated with the adapter, indicating its efficacy in capturing generic forgery traces.
Limitations. Our method is designed specifically for face forgery detection, rather than whole face image synthesis detection. Thus it only shows decent performance on recent GAN or Diffusion model generated faces. In future work, we may further excavate the potential of CLIP and develop a universal detector capable of addressing both face forgery and whole face image synthesis detection.
5 Conclusion
In this paper, we introduce Forensics Adapter, an adapter network designed to transform CLIP into a generalizable face forgery detector. Different from existing methods, our adapter is specifically tailored to capture the unique blending boundaries characteristic of forged faces using task-specific objectives. We then introduce an interactive strategy that enables CLIP and the adapter to collaboratively focus on face forgery traces. Notably, our solution of adapting CLIP for face forgery detection is novel, addressing a gap that previous methods have not explored. Extensive experiments demonstrate that, with only trainable parameters, our adapter yields substantial performance gains, improving the performance by approximately on average across five standard datasets. We believe our method can serve as a baseline for future CLIP-based face forgery detection techniques.
Acknowledgement. This work is supported in part by the National Natural Science Foundation of China (No.62402464), Shandong Provincial Natural Science Foundation (No.ZR2024QF035), and China Postdoctoral Science Foundation (No.2021TQ0314; No.2021M703036).